OpenStudy (anonymous):

FAN AND MEDAL FOR ANSWER! What is the slope of the line passing through the points (–3, 4) and (2, –1)?

OpenStudy (nirmalnema):

y1-y1/x2-x1 4+1/-3-2 5/-5 =-1

OpenStudy (anonymous):

1.10.) x1 = - 2 y1 = 3 x2 = - 5 y2 = 6 Slope, m = (y2 - y1) / (x2 - x1) m = (6 - 3) / [- 5 - (- 2)] m = 3 / (- 5 + 2) m = 3 / - 3 m = - 1 ¯¯¯¯¯¯ 11.) 2x = 1/2 y + 4 1/2 y = 2x - 4 y = 4x - 8 m = 4 The slope of a perpendicular line will be the negative reciprocal of the slope of the original line, in this case, - 1/4. m = - 1/4 ¯¯¯¯¯¯¯¯ 12.) 1/3 x + 2/3 y = 2 2/3 y = - 1/3 x + 2 y = - 1/2 + 3 The slope of a parrallel line will have the same slope as the original line. m = - 1/2

OpenStudy (igreen):

To find the slope between two points, you have to use the formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\). In this case: \(y_2 = -1\) \(y_1 = 4\) \(x_2 = 2\) \(x_1 = -3\) Plug them in: \(m = \dfrac{-1-4}{2--3}\) Simplify: \(m = \dfrac{-1-4}{2--3}\) \(m = \dfrac{-5}{5}\) \(m = -1\)